OpenStudy (alibaby):

Solve x2 = –169. undefined 13 ±13 −13

OpenStudy (alibaby):

@GracyGirl

OpenStudy (aveline):

Have you learned about 'i'?

OpenStudy (aveline):

If not, the answer is undefined, two of the same numbers multiplied together can't be negative.

OpenStudy (sshayer):

\[x^2=169=169 \times-1=169 \times \iota ^2,x=\pm 13 \iota\]

OpenStudy (sshayer):

iota is an imaginary number.